OpenStudy (anonymous):

Hey just came across this 6 mark question , help please?? (b) Solve, for 0 ≤ θ < 360°, the equation 2 tan2 θ + sec θ = 1, giving your answers to 1 decimal place.

OpenStudy (turingtest):

is that tangent squared?

OpenStudy (anonymous):

Using sin^2x+cos^2x = 1 identity seems like a good start?

OpenStudy (anonymous):

\[\sec \theta = 1 - 2\tan2\theta\] square both sides write sec^2 as 1 + tan^2 we get 4[\tan2\theta ^{2} = \tan ^{2}\] sq root it 2 tan2(theta)=tantheta

OpenStudy (anonymous):

2sin^2 x + cosx = cos^2x This should result in a quadratic equation of cos x. Then, solve?

OpenStudy (anonymous):

tan2theta = (2tan theta)/( 1+ tan^2)

OpenStudy (anonymous):

hmmmm...that seems convoluted.

OpenStudy (anonymous):

short way .. follow gt ..

OpenStudy (anonymous):

sorry messed up in the equation writer basically i converted the equation into tan(theta)

OpenStudy (turingtest):

I messed up as well. good job GT

OpenStudy (anonymous):

im getting theta = 0 or pi/3

OpenStudy (anonymous):

but hey kinzan is that squared or 2theta ??

OpenStudy (anonymous):

because i have solved for 2 theta and gt has solved taking squared

OpenStudy (anonymous):

hey guys yeah it is tan^2

OpenStudy (anonymous):

then follow gt 's answer

OpenStudy (anonymous):

I did not give the answer. Rather the approach to convert the equation into a quadratic equation of cos(theta), which when solved will lead to the answer.

OpenStudy (mertsj):

Add two to both sides first.

OpenStudy (mertsj):

\[2\tan ^{2}\theta+2+\sec \theta=\]=3

OpenStudy (mertsj):

Now factor 2 out of the first two terms:

OpenStudy (mertsj):

\[2(\tan ^{2}\theta+1)+\sec \theta=3\]

OpenStudy (mertsj):

Now substitute sec^2 for tan^2 + 1 and subtract 3 from both sides.

OpenStudy (mertsj):

\[2\sec ^{2}\theta+\sec \theta-3=0\]

OpenStudy (mertsj):

Now factor:

OpenStudy (mertsj):

\[(2\sec \theta+3)(\sec \theta-1)=0\]

OpenStudy (mertsj):

Can you get it from here?

OpenStudy (mertsj):

\[\sec \theta=-1.5 or \sec \theta=1\]

OpenStudy (mertsj):

theta = 2.3 radians or 0 radians

OpenStudy (anonymous):

@Mertsj: Thanks ! But how do you work out theta= 2.3 radians on your calculator? Do you use the rule : sec θ=1/cosθ ??

OpenStudy (mertsj):

Yes

OpenStudy (mertsj):

So cos theta = -2/3 or 1

OpenStudy (anonymous):

you mean -3/2 right??

OpenStudy (mertsj):

No. sec theta = -3/2 so cos theta = -2/3

OpenStudy (mertsj):

So type -.666666666666666 into your calculator (which is set in radian mode) and hit the inverse cosine button.

OpenStudy (mertsj):

Or if you calculator does fractions, type in -2/3 and push inverse cosine

OpenStudy (anonymous):

i got 2.3?

OpenStudy (mertsj):

theta = 2.3 radians or 0 radians

OpenStudy (anonymous):

ohh sorry in degrees mode i got 131.8 which is correct!

OpenStudy (anonymous):

yes, thanks! :D

OpenStudy (mertsj):

Both are correct. 131.8 degrees or 2.3 radians. don't forget the 0 degrees or 0 radiana

OpenStudy (anonymous):

thanks :)

OpenStudy (mertsj):

yw